OpenStudy (anonymous):

Which is the equation of a line that passes through the point (3, -4) and has a slope of 2. a. y=2x-11 b. y=2x-10 c. y=2x-4 d. y=2x-2

OpenStudy (anonymous):

is the answer c

OpenStudy (anonymous):

no it is B

OpenStudy (anonymous):

(3,-4) means that the x value is 3 and the y value is -4 2*3-10=-4

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

\[y+4=2(x-3)\] \[y+4=2x-6\] \[y=2x-10\]